OpenStudy (vortish):

use the method of substitution to solve the system of linear equations 5x+3y=5/6 y=6x-1 find the ordered pairs

OpenStudy (anonymous):

Just replace y in the first equation with the expression of the second equation

OpenStudy (vortish):

ok so it would be 5x+3(6x-1)=5/6

OpenStudy (anonymous):

Yup

OpenStudy (vortish):

ok ty

OpenStudy (asnaseer):

are you stuck on this? from the chat above it looks like you know what to do now.

OpenStudy (vortish):

yea im stuck

OpenStudy (asnaseer):

what are you having difficulty with?

OpenStudy (vortish):

trying to find the ordered pairs

OpenStudy (asnaseer):

well you got this part correct: 5x+3(6x-1)=5/6 so solve this to find a value for x. Then use: y=6x-1 to get a value for y. Then your solution will be the ordered pair given by (x, y) where you replace x and y by the numbers you found above.

OpenStudy (vortish):

well i have come up with 5x+18x=1=5/16 then 23x-1=5/16 and then i get stuck

OpenStudy (asnaseer):

so you do not know how to solve for x - looks like you first need to learn some algebra. I would recommend that you view these short videos: http://www.khanacademy.org/math/algebra/introduction-to-algebra that will give you a good grounding on alegra which will make it easier for you to solve questions like this one.

OpenStudy (vortish):

you have to get x on one side by it self

OpenStudy (asnaseer):

yes

OpenStudy (vortish):

so yo have to add like terms to do that right

OpenStudy (asnaseer):

yes, but your initial expansion of: 5x+3(6x-1)=5/6 was incorrect

OpenStudy (asnaseer):

what is: 3(6x-1)

OpenStudy (asnaseer):

when expanded?

OpenStudy (vortish):

so i come up with 23x -1=5/6 divide both the 23 and the one by 23 to come up with x-1/23=5/6 then subtract both sides by 1/23

OpenStudy (vortish):

so you come up with x = 5/6-1/23

OpenStudy (asnaseer):

please look above ^^^ you got the initial expansion wrong

OpenStudy (asnaseer):

what is: 3(6x-1) when expanded?

OpenStudy (vortish):

oh its 18x -3

OpenStudy (asnaseer):

correct, you start again from: 5x+3(6x-1)=5/6

OpenStudy (asnaseer):

also - to simplify your life - I would first get rid of the fractional parts in this equation. you can do that by multiplying both sides by 6.

OpenStudy (vortish):

5x +18x-3=5/6

OpenStudy (asnaseer):

correct - now multiply both sides by 6 to get rid of the fraction

OpenStudy (vortish):

so it would be (23x)6-3=5/6(6)

OpenStudy (asnaseer):

when you multiply both sides - you must ensure that EVERY term is multiplied

OpenStudy (asnaseer):